GAS PHASE ROVIBRATIONAL SPECTROSCOPY OF DMSO, PART II: TOWARDS THE TERAHERTZ OBSERVATION OF 4-FOLD CLUSTERS

\underline{ARNAUD CUISSET}, MARIE-ALINE MARTIN-DRUMEL, FRANCIS HINDLE, GAEL MOURET, DMITRII A. SADOVSKII , Laboratoire de Physico-Chimie de l’Atmosphère, 189A Ave. Maurice Schumann, 59140 Dunkerque, France.

Benefiting of the exceptional properties of the AILES synchrotron beamline^{a}, the gas phase Far-IR spectrum of DMSO has been recorded and resolved ^{b}. The rovibrational analysis allowed to discover a new rotational behaviour for a polyatomic molecule: the gyroscopic destabilization. ^{c} In order to explain this phenomenon, we looked for four-fold energy clusters in the high resolution ground state THz spectrum of DMSO recorded with a sub-THz spectrometer based on a frequency multiplication chain ^{d}. Pure rotational lines in the 5 lowest vibrationnally excited levels have been recorded below 700 GHz. With near 1000 rotational transitions assigned, high quantum numbers have been reached allowing to discover sequence of four-fold clusters in the out of plane bending mode of DMSO and to study the vibrational dependence of an unusual rotational dynamics.

aJ. B. Brubach et al., AIP Conf. Proc., 1214, (81), 2010.

bA. Cuisset, L. Nanobashvili, I. Smirnova, R. Bocquet, F. Hindle, G. Mouret, O. Pirali, P. Roy, D. Sadovskii,Chem. Phys. Lett., 492,(30),2010

cA. Cuisset, O. Pirali, D. Sadovskii,Phys. Rev. Lett., 109,(094101), 2012.

dG. Mouret, M. Guinet, A. Cuisset, L. Croizet, S. Eliet, R. Bocquet, F. Hindle, IEEE Sensors Journal, 13, 1, 2013.